### v1

@@ 5.1 Containment relationships of feasible regions
Talking about $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$ and $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$, the previous theorem only discusses the containment relationships for an arbitrary policy $\pi$. This policy $\pi$ may not come from the optimization of virtual-time OCP. In safe RL, besides trying to identify the maximum EFR, one also seeks to find the best policy, i.e., $\pi^{*}$. The optimal policy $\pi^{*}$, including its corresponding IFR and EFR, have special importance in building effective RL algorithms. As a special case, let us first talk about an interesting property of IFR at $\pi^{*}$, i.e., $\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})$. For any state $\forall x\in\mathrm{X}_{\mathrm{init}}^{g}$, its corresponding virtual-time OCP must has a feasible solution, and one can construct an optimal policy $\pi^{*}$ in this region by the mapping from this state to its feasible action. Therefore, $x$ must be initially feasible under the policy $\pi^{*}$, i.e., $x\in\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})$. This leads to an important condition that $\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})=\mathrm{X}_{\mathrm{init}}^{g}$, and therefore a series of useful properties.
@@
    According to Theorem 5.1(2), it holds that $\mathrm{X}_{\mathrm{edls}}^{g}(\pi^{*})\subseteq\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})\subseteq\mathrm{X}_{\mathrm{init}}^{g}$. What left is to show that $\mathrm{X}_{\mathrm{init}}^{g}\subseteq\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})$ and $\mathrm{X}_{\mathrm{edls}}^{g}\subseteq\mathrm{X}_{\mathrm{edls}}^{g}(\pi^{*})$. Since
@@
    we have $\mathrm{X}_{\mathrm{init}}^{g}\subseteq\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})$, which yields $\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})=\mathrm{X}_{\mathrm{init}}^{g}$. Now we can choose $\pi=\pi^{*}$ in Theorem 5.1(4) and conclude that $\mathrm{X}_{\mathrm{edls}}^{g}\subseteq\mathrm{X}_{\mathrm{edls}}^{g}(\pi^{*})$. Thus, we have
@@
    This follows directly from Corollary 5.1.1(1) and Theorem 5.1(3).
@@
Corollary 5.1.1(1) reveals the relationship between IFR, EFR, and those of the optimal policy. This relationship tells us a range of the optimal policy’s EFR, i.e., it is lower bounded by EFR and upper bounded by the optimal policy’s own IFR. Corollary 5.1.1(2) further clarifies the range of the optimal policy’s EFR by revealing its relationship with the maximum EFR, i.e., it is a subset of the latter. Corollary 5.1.1(3) gives a method for obtaining the EFR of the optimal policy under a certain condition that IFR equals EFR. In this case, the EFR of the optimal policy can be easily obtained since it directly equals the IFR. Therefore, this condition also provides a rule for designing virtual-time constraints. Corollary 5.1.1(4) gives a condition when the EFR of the optimal policy equals the maximum EFR. The condition is that the EFR equals the maximum EFR. Combining with Corollary 5.1.1(3), we arrive at an important rule for designing virtual-time constraints: $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}=\mathrm{X}_{\mathrm{edls}}^{*}$, i.e., EFR, IFR, and the maximum EFR are equal. Following this rule, we can guarantee that the EFR of the optimal policy equals the maximum EFR.
@@ 5.2 Equivalence conditions of two feasible regions
The EFR is only determined by the dynamics model and its virtual-time constraint. The perfect design for a virtual-time constraint should result in an EFR that equals the maximum EFR, i.e., $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{edls}}^{*}$. However, $\mathrm{X}_{\mathrm{edls}}^{g}$ is very difficult to compute because it requires checking every initially feasible policy in every time step. To avoid this problem, we hope that $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}$ so that we only need to compute $\mathrm{X}_{\mathrm{init}}^{g}$, which is much easier than computing $\mathrm{X}_{\mathrm{edls}}^{g}$. In this case, if $\mathrm{X}_{\mathrm{init}}^{g}=\mathrm{X}_{\mathrm{edls}}^{*}$, then $\mathrm{X}_{\mathrm{edls}}^{g}(\pi^{*})=\mathrm{X}_{\mathrm{edls}}^{*}$. The following theorem helps us to have a sense of what kind of condition $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}$ needs.
Theorem 5.2 (Conditions for $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}$).
@@
    This is obvious by noting Definition 4.5.
@@
Theorem 5.2 has important guiding significance for designing virtual-time constraints. The necessary conditions should always be satisfied in constraint design. These conditions require that IFR is a subset of the maximum EFR and a control invariant set at the same time. Checking these conditions only involves computing IFR, which is much easier than computing EFR. One may argue that the maximum EFR is unknown so the first necessary condition cannot be checked. In fact, we do not need to exactly know the maximum EFR. Instead, we just need to check if all states in IFR are endlessly feasible for a given policy. If this is true, then the first necessary condition is satisfied. The sufficient conditions in Theorem 5.2 can be viewed as two examples of virtual-time constraints that can satisfy $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}$. When designing virtual-time constraints, one should try to satisfy a sufficient condition under the premise that the necessary conditions are already satisfied. Note that there still exist other forms of sufficient conditions that can ensure the equivalence of EFR and IFR. We can choose different ones according to the specific constrained OCP we aim to solve.
6 Feasibility Function and Constraint Types
The goal of safe RL is to find not only the optimal policy but also its EFR. Whether this EFR equals the maximum EFR depends on the choice of virtual-time constraints. Therefore, constructing a proper virtual-time constraint becomes an indispensable task in constrained optimal control. In this section, we first introduce a tool for representing EFR called feasibility function, which helps us construct proper constraints that satisfy the equivalence conditions and enable us to find the maximum EFR. Next, we review several commonly used virtual-time constraint formulations from the perspective of feasibility function. We point out that these virtual-time constraints are all constructed from some feasibility function and therefore satisfy some desirable properties. Moreover, we use the tools introduced in the previous section to analyze the containment relationships of feasible regions under these virtual-time constraints.
@@ 6.1 Feasibility function
Feasibility functions are used for constructing virtual-time constraints and representing EFRs. A feasibility function is defined as a mapping from the state space to a real number, i.e., $F:\mathcal{X}\to\mathbb{R}$. Through proper design, its zero-sublevel set can represent an EFR:
$\mathrm{X}_{F}=\{x|F(x)\leq 0\}.$
The core of designing feasibility functions is to offer some kind of recursion properties so that by properly constructing the virtual-time constraint, $\mathrm{X}_{F}$ can become an EFR. There are two basic methods that lead to the design of feasibility functions. The first is defined through a control invariant set (CIS) and the second is defined through constraint aggregation (CA). A control invariant set is a region in which there exists a policy that keeps all successive states still in this region. A constraint aggregation is a function that can replace the infinite-step real-time constraint with a single-step virtual-time constraint. These two kinds of feasibility functions result in two families of virtual-time constraints. The former type restricts the state in a control invariant set while the latter type constrains an aggregation of infinite-step real-time constraints.
For the first type, the zero-sublevel set of feasibility function is chosen as a control invariant set, i.e., the state is maintained within this set under certain control given that the previous state is in the set. This type of feasibility function is defined as follows.
Definition 6.1 (Control invariant set).
@@
$\forall x\in\mathrm{X}_{F},\exists u,\text{ s.t. }x^{\prime}\in\mathrm{X}_{F}.$
@@
In general, certain inequality conditions are constructed with $F$ to serve as virtual-time constraints when using this type of feasibility function. These constraints require the policy to yield a next state that is still in $\mathrm{X}_{F}$, starting from whatever state in $\mathrm{X}_{F}$. That is to say, through the virtual-time constraint built upon it, the feasibility function $F$ equips the resulting policy with the recursive property we are looking for. But does this always work? In other words, how can we be sure that such a policy that satisfies the constraint and thus has the recursive property always exists? The answer is through the special property of the zero-sublevel set of $F$, i.e., control invariance. Figure 4 gives an illustration of a feasibility function and its control invariant set. Any state $x$ inside the control invariant set can be still kept in this set under some action $u$, while a state $\tilde{x}$ outside the control invariant set may not be able to enter this set.
Figure 4: Feasibility function defined through control invariant set.
@@
For the second type, the feasibility function is chosen as an aggregation function whose one-step virtual-time constraint can represent the infinite-step real-time constraint. This type of feasibility function is defined as follows.
Definition 6.2 (Constraint aggregation).
A function $F:\mathcal{X}\to\mathbb{R}$ is a feasibility function if $\forall x\in\mathcal{X}$,
$F(x)\leq 0\iff h(x_{t+i})\leq 0,x_{t}=x,i=0,1,\dots,\infty.$
@@
Figure 5 gives an illustration of a feasibility function defined through constraint aggregation. For states $x_{1}$ and $x_{3}$, although they do not violate the state constraint at the current step, some states on their future trajectories leave the constrained set. Therefore, the feasibility function $F$ is positive on these two states. For state $x_{2}$, all of its successive states are inside the constrained set, making the feasibility function less than or equal to zero. Definition 6.2 means that $F$ must be a function of a policy $\pi$. We can also denote it as $F^{\pi}$ to show its connection with $\pi$. This equivalence compresses the infinitely many constraints into $F$ and hence equips it with a natural recursive property, i.e., once $F(x_{t})\leq 0$, it follows that $F(x_{t+1})\leq 0$. However, this recursive property is not a free lunch. We have to make sure that the value of $F$ is practically available before it can be used to construct virtual-time constraints. One may recall that the state-value function and the action-value function in RL also contain infinite future rewards, and we can calculate them iteratively by bootstrapping. Similarly, feasibility functions of constraint aggregation type are usually formulated in such ways that they satisfy a self-consistency condition and hence can be calculated iteratively.
Figure 5: Feasibility function defined through constraint aggregation.

### v2

@@ 5.1 Containment relationships of feasible regions
Talking about $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$ and $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$, the previous theorem only discusses the containment relationships for an arbitrary policy $\pi$. This policy $\pi$ may not be the solution to a virtual-time OCP. In safe RL, besides trying to identify the maximum EFR, one also seeks to find the optimal policy, i.e., $\pi^{*}$. Let us start from the IFR of $\pi^{*}$, i.e., $\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})$. For any state $x\in\mathrm{X}_{\mathrm{init}}^{g}$, its corresponding virtual-time OCP must have feasible solutions, and $\pi^{*}$ is one of them. Therefore, $x$ must be initially feasible under $\pi^{*}$, i.e., $x\in\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})$. This leads to an important condition $\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})=\mathrm{X}_{\mathrm{init}}^{g}$, and therefore a series of useful properties.
@@
    According to Theorem 5.1.1(2), it holds that $\mathrm{X}_{\mathrm{edls}}^{g}(\pi^{*})\subseteq\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})\subseteq\mathrm{X}_{\mathrm{init}}^{g}$. What left is to show that $\mathrm{X}_{\mathrm{init}}^{g}\subseteq\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})$ and $\mathrm{X}_{\mathrm{edls}}^{g}\subseteq\mathrm{X}_{\mathrm{edls}}^{g}(\pi^{*})$. Since
@@
    we have $\mathrm{X}_{\mathrm{init}}^{g}\subseteq\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})$, which yields $\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})=\mathrm{X}_{\mathrm{init}}^{g}$. Now we can choose $\pi=\pi^{*}$ in Theorem 5.1.1(4) and conclude that $\mathrm{X}_{\mathrm{edls}}^{g}\subseteq\mathrm{X}_{\mathrm{edls}}^{g}(\pi^{*})$. Thus, we have
@@
    This follows directly from Corollary 5.1.3(1) and Theorem 5.1.1(3).
@@
Corollary 5.1.3(1) reveals that the optimal policy’s EFR is lower bounded by EFR and upper bounded by its own IFR. Corollary 5.1.3(2) further clarifies that the optimal policy’s EFR is a subset of the maximum EFR. Corollary 5.1.3(3) gives a method for obtaining the optimal policy’s EFR under a certain condition that IFR equals EFR. This condition also serves as a rule for designing virtual-time constraints. Corollary 5.1.3(4) gives a condition when the optimal policy’s EFR equals the maximum EFR: the EFR equals the maximum EFR. Combining with Corollary 5.1.3(3), we arrive at an important rule for designing virtual-time constraints: $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}=\mathrm{X}_{\mathrm{edls}}^{*}$, i.e., EFR, IFR, and the maximum EFR are equal. Following this rule, we can guarantee that the optimal policy’s EFR equals the maximum EFR.
The sizes of feasible regions are largely determined by virtual constraints. As we mentioned in Chapter 3, weaker virtual-time constraints have worse feasibility in the long run. Here, we formally state and prove how the strength of virtual-time constraints affects the sizes of feasible regions. We begin with the following lemma.
Lemma 5.1.5.
For all $x\in\mathrm{X}_{\mathrm{edls}}^{g}$ and $\pi\in\Pi_{\text{init}}^{g}(x)$, $x_{t+1}\in\mathrm{X}_{\mathrm{edls}}^{g}$.
Proof 5.1.6.
This holds from the arbitrariness of policies and the infinity in Definition 4.2.1.
The following theorem states that a weaker virtual-time constraint leads to a smaller EFR.
Theorem 5.1.7.
Given two virtual-time constraints $\text{VC}_{1}$ and $\text{VC}_{2}$, we term $\text{VC}_{1}$ is weaker than $\text{VC}_{2}$ if starting from any state $x$, a policy satisfying $\text{VC}_{2}$ also satisfies $\text{VC}_{1}$. Let $\text{VC}_{h}$ be a special virtual-time constraint that is exactly the same with the real-time constraint. Suppose that $\text{VC}_{1}$ is weaker than $\text{VC}_{2}$, and $\text{VC}_{2}$ is weaker than $\text{VC}_{h}$, then $\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{1}}\subseteq\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{2}}\subseteq\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{h}}=\mathrm{X}_{\mathrm{edls}}^{*}=\mathrm{X}_{\mathrm{init}}^{\text{VC}_{h}}\subseteq\mathrm{X}_{\mathrm{init}}^{\text{VC}_{2}}\subseteq\mathrm{X}_{\mathrm{init}}^{\text{VC}_{1}}$.
Proof 5.1.8.
It holds by Definition 4.1.1, Definition 4.2.1 and Theorem 4.2.4 that $\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{h}}=\mathrm{X}_{\mathrm{edls}}^{*}=\mathrm{X}_{\mathrm{init}}^{\text{VC}_{h}}\subseteq\mathrm{X}_{\mathrm{init}}^{\text{VC}_{2}}\subseteq\mathrm{X}_{\mathrm{init}}^{\text{VC}_{1}}$. The left is to show that $\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{1}}\subseteq\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{2}}\subseteq\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{h}}$. We only proof the first containment and the second one follows the same procedure. For all $x_{t}\in\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{1}}$, by Definition 4.2.1 and Lemma 5.1.5, we have $\forall\pi_{t+i}\in\Pi_{\mathrm{init}}^{\text{VC}_{1}}(x_{t+i}),$ $x_{t+i+1}\in\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{1}},i=0,1,2,\dots,\infty$. Together with the facts that $\Pi_{\mathrm{init}}^{\text{VC}_{2}}(x_{t+i})\subseteq\Pi_{\mathrm{init}}^{\text{VC}_{1}}(x_{t+i})$ and that $\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{1}}\subseteq\mathrm{X}_{\mathrm{edls}}^{*}\subseteq\mathrm{X}_{\mathrm{init}}^{\text{VC}_{2}}$, we can conclude that $x_{t}\in\mathrm{X}_{\mathrm{init}}^{\text{VC}_{2}}$ and $\forall\pi_{t+i}\in\Pi_{\mathrm{init}}^{\text{VC}_{2}}(x_{t+i}),$ $x_{t+i+1}\in\mathrm{X}_{\mathrm{init}}^{\text{VC}_{2}},i=0,1,2,\dots,\infty$, which means $x_{t}\in\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{2}}$. Thus, $\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{1}}\subseteq\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{2}}$.
Remark 5.1.9.
It is worth distinguishing the case where virtual-time constraints are tighter than the real-time constraint $\text{VC}_{h}$. Suppose that $\text{VC}_{h}$ is weaker than $\text{VC}_{1}$, and $\text{VC}_{1}$ is weaker than $\text{VC}_{2}$ (i.e., $\text{VC}_{2}$ is the strictest). In this case, although $\mathrm{X}_{\mathrm{init}}^{\text{VC}_{2}}\subseteq\mathrm{X}_{\mathrm{init}}^{\text{VC}_{1}}$, the relationship $\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{2}}\subseteq\mathrm{X}_{\mathrm{edls}}^{\text{VC}_{1}}$ does not necessarily hold. This is because $\mathrm{X}_{\mathrm{edls}}$ requires the safety condition to hold for all initially feasible policies. Since $\Pi_{\mathrm{init}}^{\text{VC}_{2}}(x)\subseteq\Pi_{\mathrm{init}}^{\text{VC}_{1}}(x)$, the fact that all policies in the smaller set $\Pi_{\mathrm{init}}^{\text{VC}_{2}}(x)$ maintain feasibility does not guarantee that the additional policies in $\Pi_{\mathrm{init}}^{\text{VC}_{1}}(x)$ also do so. Therefore, satisfaction of the endless feasibility condition under $\text{VC}_{2}$ does not imply satisfaction under $\text{VC}_{1}$.
@@ 5.2 Equivalence conditions of two feasible regions
The perfect design for a virtual-time constraint should result in an EFR that equals the maximum EFR, i.e., $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{edls}}^{*}$. However, $\mathrm{X}_{\mathrm{edls}}^{g}$ is difficult to compute because it requires checking every initially feasible policy in every time step. To avoid this problem, we hope that $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}$ so that we only need to compute $\mathrm{X}_{\mathrm{init}}^{g}$, which is much easier than computing $\mathrm{X}_{\mathrm{edls}}^{g}$. In this case, if $\mathrm{X}_{\mathrm{init}}^{g}=\mathrm{X}_{\mathrm{edls}}^{*}$, then $\mathrm{X}_{\mathrm{edls}}^{g}(\pi^{*})=\mathrm{X}_{\mathrm{edls}}^{*}$. The following theorem helps us understand what kind of condition $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}$ needs.
Theorem 5.2.1 (Conditions for $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}$).
@@
    This is obvious by noting Definition 4.2.3.
@@
Theorem 5.2.1 has important guiding significance for designing virtual-time constraints. One may argue that since the maximum EFR is unknown, the first necessary condition cannot be checked. In fact, we do not need to know the maximum EFR exactly. Instead, we can check if all states in IFR are endlessly feasible for a given policy. If this is true, the first necessary condition is satisfied. However, it remains intractable to check endless feasibility by definition in high-dimensional spaces, since we cannot traverse the state space. Instead, we can only prove or falsify endless feasibility by theoretical analysis. As we will demonstrate in Chapter 7, this is achievable for established feasibility functions, such as CBF and Hamilton-Jacobi reachability function. The sufficient conditions in Theorem 5.2.1 can be viewed as two examples of virtual-time constraints satisfying $\mathrm{X}_{\mathrm{edls}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}$. When designing virtual-time constraints, one should try to satisfy a sufficient condition under the premise that the necessary conditions are satisfied. Note that other forms of sufficient conditions still exist that can ensure the equivalence of EFR and IFR. We can choose different ones according to the specific constrained OCP we aim to solve.
Chapter 6 Feasibility Function and Constraint Types
Solving constrained OCPs involves finding both the optimal policy and its EFR. Whether the EFR equals the maximum EFR depends on the choice of virtual-time constraints. Therefore, constructing a proper virtual-time constraint is an indispensable task in constrained optimal control. In this chapter, we first introduce a tool for representing EFR called feasibility function, which helps us construct proper constraints that satisfy the equivalence conditions and enable us to find the maximum EFR. Next, we review several commonly used virtual-time constraint formulations from the perspective of feasibility function. We point out that these virtual-time constraints are all constructed from feasibility functions and, therefore, satisfy desirable properties. Moreover, we use the tools introduced in the previous chapter to analyze the containment relationships of feasible regions under these virtual-time constraints.
@@ 6.1 Feasibility function
Feasibility functions are used for constructing virtual-time constraints and representing EFRs. A feasibility function is a mapping from the state space to a real number, i.e., $F:\mathcal{X}\to\mathbb{R}$. Through proper design, its zero-sublevel set $\mathrm{X}_{F}=\{x|F(x)\leq 0\}$ represents an EFR. There are two basic methods for designing feasibility functions. The first uses a control invariant set and the second resorts to constraint aggregation. A control invariant set is a region where there exists a policy that keeps all successive states still in this region. A constraint aggregation equivalently replaces the infinite-step real-time constraint with a single-step virtual-time constraint. These two types of feasibility functions result in two families of virtual-time constraints.
For the first type, the zero-sublevel set of the feasibility function is chosen as a control invariant set, which is defined as follows.
Definition 6.1.1 (Control invariant set (CIS)).
@@
[…]
where $x^{\prime}=f(x,u)$.
@@
When using this type of feasibility function, certain inequalities are used as virtual-time constraints, which require the next state to be kept in $\mathrm{X}_{F}$ starting from any state in $\mathrm{X}_{F}$. Figure 6.1 gives an illustration of a feasibility function and its control invariant set. Any state $x$ inside the control invariant set can be still kept in this set under some action $u$, while a state $\tilde{x}$ outside the control invariant set may not be able to enter this set.
Figure 6.1: Feasibility function defined through control invariant set.
@@
For the second type, the feasibility function is chosen as an aggregation function of the real-time constraint, which is defined as follows.
Definition 6.1.2 (Constraint aggregation (CA)).
A function $F:\mathcal{X}\to\mathbb{R}$ is a feasibility function if
$\exists\pi,\text{ s.t. }\forall x\in\mathcal{X},F(x)\leq 0\iff h(x_{t})\leq 0,t=0,1,\dots,\infty,$
where $x_{0}=x$, $x_{t+1}=f(x_{t},\pi(x_{t}))$.
@@
Figure 6.2 gives an illustration of constraint aggregation. For states $x_{1}$ and $x_{3}$, although they do not violate the state constraint at the current step, their future states leave the constrained set. Therefore, the feasibility function $F$ is positive on these two states. For state $x_{2}$, all of its successive states are inside the constrained set, thus its feasibility function is less than or equal to zero. Definition 6.1.2 implies that $F$ is a function of a policy $\pi$. We can also denote it as $F^{\pi}$ to show its connection with $\pi$. One may recall that the state-value and action-value functions in RL also contain infinite future rewards, and we can compute them iteratively by bootstrapping. Similarly, feasibility functions of this type are usually formulated to satisfy self-consistency conditions so that they can be computed iteratively.
The control invariant set method is most suitable for low-dimensional systems with known dynamics, as constructing such sets typically requires explicit knowledge of the system dynamics and remains computationally tractable only in lower dimensions. This approach provides strict, provable safety guarantees, but it may sacrifice optimality—especially when the resulting invariant set is overly conservative. It is thus well-suited for applications where safety is critical and some degree of optimality can be compromised.
In contrast, the constraint aggregation method can be applied to higher-dimensional systems and those with unknown dynamics. Feasibility functions designed via constraint aggregation generally satisfy self-consistency conditions, enabling them to be solved via fixed-point iteration. When parameterized using neural networks, such functions can be optimized iteratively via dynamic programming, much like value functions in reinforcement learning. A key advantage of this approach is its theoretical ability to recover the maximum feasible region in any system. However, this scalability comes at a cost: the zero-sublevel sets of these feasibility functions are not necessarily forward invariant, which implies that strong safety guarantees may not always be available. This behavior mainly stems from approximation and generalization errors when feasibility functions are represented by neural networks. Imperfect training, limited data, and finite network capacity prevent exact recovery of the true feasibility function.
The feasibility function has a natural connection to the value function defining the infinite-horizon backward avoidable tube in Hamilton-Jacobi reachability analysis. This reachability value function, whose zero-sublevel set is the maximum EFR, is a quintessential example of a feasibility function. However, the concept of a feasibility function is more general. Its zero-sublevel set is not restricted to the maximum EFR but can represent the infinite-horizon constraint-satisfying set of any policy. Moreover, its functional form is not confined to the specific “maximum-over-time” structure of the reachability solution. As we will show in the following two sections, the Hamilton-Jacobi value function is a specific instance within a broader and more flexible class of the feasibility function.
Figure 6.2: Feasibility function defined through constraint aggregation.